# Protected gap closing and reopening in topological-insulator Josephson junctions

### v1

@@ Abstract
In the seminal proposal by Fu and Kane, the superconducting proximity effect was predicted to transform the surface state of a topological insulator (TI) into a topological superconductor, forming a nonchiral 1D Majorana state within a linear Josephson junction on the TI surface. The hallmark of this 1D Majorana state is a robust gap closing as a function of the superconducting phase difference $\varphi$ across the junction, which alternates in and out of the topological phase. These topological phase-transitions occur at $\varphi=(2n+1)\pi$ with integer $n$, leading to a $4\pi$-periodicity of the ground state. While the $4\pi$-periodicity has been indirectly inferred in the AC Josephson effect, the direct observation of the 1D Majorana state in a TI Josephson junction has remained contentious. Here, we report the direct observation of topological phase-transitions in a TI Josephson junction, where the local density of states is probed via tunnel contacts and $\varphi$ is controlled by a flux loop. The observed transitions are independent of the chemical potential, reinforcing their topological origin. Under an applied perpendicular magnetic field, Josephson vortices form, making $\varphi$ position-dependent. In this case, the gap closing occurs locally at the Josephson vortex cores where $\varphi=(2n+1)\pi$, which we also observe. Our findings provide direct confirmation of the Fu-Kane proposal and robust evidence for the emergence of topological superconductivity in a TI Josephson junction.
{affiliations}
Physics Institute II, University of Cologne, Zülpicher Str. 77, 50937 Köln, Germany
Institute for Quantum Information, RWTH Aachen University, 52056 Aachen, Germany
Max Planck Institute for the Science of Light, 91058 Erlangen, Germany
Department of Physics, Friedrich-Alexander-Universität Erlangen-Nürnberg, 91058 Erlangen, Germany
These authors contributed equally
1 Introduction
@@
After more than 100 years from the discovery, superconductors remain a fascinating subject of research in physics. They become even more interesting when topology comes into play[1]. A special type of superconductor, called “spinless chiral $p$-wave superconductor”, is a prototypical topological superconductor hosting Majorana zero-modes at point-defects[2, 1]. The Majorana zero-modes obey non-Abelian exchange statistics[3, 4], which make them a promising candidate for topological quantum computing[5]. Due to this exciting prospect for applications, the realization of topological superconductivity and the observation of Majorana zero-modes have become major themes in condensed matter physics.
@@
Since a spinless chiral $p$-wave superconductor is not yet found in nature, there have been significant efforts to effectively realize it in a hybrid system[6, 7]. Theoretically, a hybrid of a strongly spin-orbit-coupled material and a conventional superconductor would effectively realize the spinless chiral $p$-wave pairing due to spin-momentum locking. The earliest of such proposals made by Fu and Kane in 2008 is based on topological insulator (TI) surface states proximitised by a conventional $s$-wave superconductor[8]. However, even though the spin-momentum locking is most naturally realized in TIs, the experimental evidence for the realization of topological superconductivity obtained so far in this platform is weak and relies on indirect signatures, such as modified Fraunhofer patterns[9, 10, 11] and the microwave response[12, 13, 14] of a Josephson junction made on a TI, or a broad zero-bias conductance peak (ZBCP) observed in the Abrikosov vortex core of a proximitised TI[15]. The search for Majorana zero-modes has been more advanced on one-dimensional (1D) semiconducting Rashba nanowires, proximity-coupled to a superconductor, and exposed to a strong magnetic field[7]; however, despite a remarkable experimental progress[16, 17], the distinction between topological and trivial ZBCPs remains contentious in this 1D platform[18, 19, 20]. It has been realized early on[21] and emphasized in the recent literature[22, 23, 24, 25] that the direct observation of the topological phase-transition, the closing and reopening of the superconducting (SC) gap, is a more robust signature of a topological superconductor and thus Majorana zero-modes.
In this work, we turn to the TI platform and report the observation of protected gap closing and reopening in a TI Josephson junction upon changing the phase difference $\varphi$ across the junction, the key signature of the 1D Majorana state predicted by Fu and Kane for a phase-biased TI Josephson junction realizing 1D topological superconductivity[8]. The gap closing and reopening correspond to a switch in the ground-state parity, which implies a change in the $\mathbb{Z}_{2}$ topological invariant [2] and signals a topological phase-transition of the 1D Majorana state formed in the junction. The phase transition is topologically-protected in the absence of parity-mixing terms. At the phase transition, which occurs at $\varphi=(2n+1)\pi$ with integer $n$, the gap closes and a gapless 1D mode appears. Note that the Fu-Kane theory is only valid for a constant $\varphi$ along the junction. When a perpendicular magnetic field is applied to the junction, rendering $\varphi$ position-dependent[26], the gapless 1D mode becomes confined at the core of a Josephson vortex, where the local phase difference is $\varphi=(2n+1)\pi$. Because of this, the gapless 1D mode evolves into a Majorana zero-mode that is bound to the point defect formed by the Josephson vortex. In our experiment, we observe a periodic closing and reopening of the gap by tuning the phase difference across the junction in a SQUID geometry and simultaneously detecting the local density of states (LDOS) at different locations. The gap closing is accompanied by a multitude of low-lying excited states, as expected from theory. We further elucidate the locality of the involved states created by the phase gradient in perpendicular magnetic fields. Our observation is a direct experimental proof of the Fu-Kane theory and is a strong indication of the emergence of topological superconductivity in a TI Josephson junction. Based on this experiment, we discuss the desirable parameter range to realize well-protected Majorana zero-modes in a TI Josephson junction, which turns out to be within experimental reach.
2 Phase-biased topological-insulator Josephon junction
@@
The 1D Majorana state predicted by Fu and Kane has a $4\pi$-periodic $\varphi$-dependence at zero momentum as shown in Fig. 1a. The corresponding evolution of the gap in the junction is depicted in Fig. 1c: The gap is trivial for $0\leq\varphi<\pi$. It closes at the topological phase-transition with $\varphi=\pi$ where the gapless 1D mode (green line) appears, before reopening in the topological regime, with $\pi<\varphi<3\pi$. In the topological regime the parity of the ground state is odd, meaning that the bound state below the chemical potential $\mu$ is filled with an unpaired electron (shown in yellow). Because of this, the junction returns to the initial state only at $\varphi=4\pi$. In contrast, for a conventional Andreev bound state, a $2\pi$-periodicity is expected, as depicted in Fig. 1b. Here, the dispersion is given by[27]
@@
with $\tau$ the transmittance of the state and $\Delta$ the induced SC gap in the TI surface directly beneath the SC electrodes. In such a scenario, the gap can only close for the fine-tuned case of perfect transmission ($\tau$ = 1). Moreover, in this case, the ground-state parity would remain even across the gap closing as a conventional Andreev bound state is doubly degenerate due to the spin degeneracy.
@@
A scanning electron microscope (SEM) image of our typical device is shown in Fig. 1d. A Josephson junction is created by a thin film of Al on top of an exfoliated flake of the bulk-insulating TI material BiSbTeSe2[28, 29]. The two sides of the junction are connected by an Al loop, which allows us to accurately tune the phase difference $\varphi$ by applying a perpendicular magnetic field. Such a device geometry is commonly used to study the bound state spectrum in an SNS-type Josephson junction via tunnel spectroscopy[30, 31, 32, 33]. Normal electrodes made of Au are deposited after putting a thin layer of $\mathrm{Al_{2}O_{3}}$ as a tunnel barrier. Due to the native oxide of Al, its surface can be treated as an insulator and the overlap to the Au electrodes does not contribute to the transport. A back-gate voltage can be applied through the degenerately-doped silicon substrate. Therefore, this setup enables us to directly probe the LDOS of the TI Josephson junction and to test the Fu-Kane theory[8]. Note that when the perpendicular magnetic field is very weak such that the large loop is threaded by only a few flux quanta (and no Josephson vortex nucleates in the junction), the phase $\varphi$ is almost constant along the junction, a situation akin to the Fu-Kane model. The case at larger magnetic field, when Josephson vortices nucleate in the junction, is discussed later. We focus on the results obtained from a device similar to the one shown in Fig. 1d with two independent tunnel junctions TJ1 and TJ2 located at different positions on the same Josephson junction. Additional data obtained from a second, essentially identical device, is shown in the supplement to demonstrate the reproducibility of our results.
3 Tunnel spectra and protected gap closing
@@
In Fig. 1e, we show the continuous evolution of the differential conductance $dI/dV$ of TJ1 with the applied DC bias voltage $V_{b}$ and the perpendicular magnetic field $B$ around zero. Line cuts for selected fields are shown in Fig. 1f. Without magnetic field applied, we observe sharp particle-hole symmetric peaks at the bias voltages of ±$\pm 130\,\mathrm{\mu eV}$ and a strong reduction in $dI/dV$ at lower bias. The constant $dI/dV$ at higher bias corresponds to a resistance of about $400\,\mathrm{k\Omega}$, indicating we are in the tunnelling regime, where the observed $dI/dV$ is proportional to the LDOS. The peak energy of $130\,\mathrm{\mu eV}$ is smaller than the SC gap of our Al-film, which can be calculated from the critical temperature to be $170\,\mathrm{\mu eV}$. This confirms that the LDOS of the proximitised TI surface, not the LDOS of the Al electrode, is being measured. The LDOS at low bias does not drop to zero but remains finite, which indicates additional in-gap states that are not fully proximitised, consistent with our previous result on similar TI Josephson junctions[34]. Such a soft gap has also been reported for comparable devices based on carbon nanotube[30], graphene[31], HgTe[32], or InAs[33].
@@
As a finite phase bias $\varphi$ is generated across the junction with the application of a perpendicular magnetic field, the positions of the particle-hole symmetric peaks shift to a lower bias voltage and their amplitude is diminished. For the magnetic field of about $115\,\mathrm{\mu T}$ that corresponds to $\varphi=\pi$, the particle-hole symmetric peaks merge into a single broad peak at zero bias, signalling the closing of the gap. The spectrum then evolves periodically, with the conductance at $230\,\mathrm{\mu T}$ becoming identical to the one observed at zero field. Having the Fu-Kane theory in mind, our data is consistent with the interpretation that the peak positions trace the gap, corresponding to the lowest-lying state at zero momentum, that evolves with $\varphi$ and closes at $\pi$ (see Fig. 1a). Above this gap (but still below the bulk gap), there is a continuum of states corresponding to the motion along the junction[8]. Our LDOS spectra reflects this physics showing a higher conductance above the gap when compared to the value below. At $\varphi=\pi$, the gap closing indicates a topological phase-transition[8].
It is important to note that the gap closing is observed despite the moderate interface transparency of our junction. Moderate transparency indicates that the junction transmittance $\tau$ is less than 1, which implies that conventional Andreev bound states should be gapped according to Eq. (1). Indeed, similar devices made on various materials reported in the literature[30, 31, 32, 33] all had $\tau<1$ and exhibited gapped spectra at $\varphi=\pi$. In contrast, our observation shows a gap closing, which is distinctly different from these results in non-topological materials supporting that the gap closing is of topological origin.
@@
To further corroborate this interpretation, we have performed numerical simulations of the differential conductance $G$ expected from the 1D Majorana state for the geometry and parameters of our devices (see supplement for details). Figure 2a shows the calculated $G$ as a function of $V_{b}$ and $\varphi$, while the line cuts for the same $\varphi$ values as in Fig. 1f are shown in Fig. 2b. These simulations clearly identify the two branches that cross at $\varphi=(2n+1)\pi$ to be the even- and odd-parity bound states having a $4\pi$-periodicity. In Fig. 2b, the coherence peaks show up at the edge of the excitation gap and they diminish rapidly with increasing $\varphi$. Most importantly, the simulations show that the gap closing at $\varphi=\pi$ results in a broad hump rather than a sharp ZBCP. All these results are consistent with our experimental observation.
4 Locality of the gap closing in perpendicular magnetic fields

### v2

# thanks: These authors contributed equally to this workthanks: These authors contributed equally to this work

@@ Abstract
In the seminal proposal by Fu and Kane, the superconducting proximity effect is used to realize topological superconductivity in the topological surface state (TSS) of a 3D topological insulator (TI). In a line Josephson junction made on the TI surface, the spin-momentum locking of the TSS guarantees the existence of a pair of spin-non-degenerate, perfectly transmitted Andreev modes. These modes lead to robust gap closing and parity alteration as a function of the superconducting phase difference $\varphi$ across the junction. Here, we report the observation of the predicted gap closing at $\varphi=(2n+1)\pi$ in a TI Josephson junction ($n$ integer), where the local density of states is probed via tunnel contacts and $\varphi$ is controlled by a flux loop. This phenomenon is robust for a wide range of chemical potentials, supporting its TSS origin. Under an applied perpendicular magnetic field, Josephson vortices form, making $\varphi$ position-dependent. In this case, the gap closing occurs locally at the Josephson vortex cores where $\varphi=(2n+1)\pi$, which we also observe. Our results confirm the fundamental role of spin-momentum locking in the Andreev physics in the TSS, which implies that the observed gap closing and reopening has a topological nature.
I Introduction
@@
After more than 100 years from the discovery, superconductors remain a fascinating subject in physics. They become even more interesting when topology comes into play [1]. A special type of superconductor, called “spinless chiral $p$-wave superconductor”, is a prototypical topological superconductor hosting Majorana zero-modes at point-defects [2, 1]. The Majorana zero-modes obey non-Abelian exchange statistics [3, 4], which make them a promising candidate for topological quantum computing [5]. Due to this exciting prospect for applications, the realization of topological superconductivity and the observation of Majorana zero-modes have become major themes in condensed matter physics.
@@
Since a spinless chiral $p$-wave superconductor is not yet found in nature, there have been significant efforts to effectively realize it in a hybrid system [6, 7]. Theoretically, a hybrid of a strongly spin-orbit-coupled material and a conventional superconductor would effectively realize the spinless chiral $p$-wave pairing due to spin-momentum locking. The earliest of such proposals made by Fu and Kane in 2008 is based on the topological surface state (TSS) of a 3D topological insulator (TI) proximitised by a conventional $s$-wave superconductor [8]. However, even though the spin-momentum locking is most naturally realized in TIs, the experimental evidence for topological superconductivity obtained so far in this platform is weak and relies on indirect signatures, such as modified Fraunhofer patterns [9, 10, 11] and the microwave response [12, 13, 14] of a Josephson junction made on a TI, or a broad zero-bias conductance peak (ZBCP) observed in the Abrikosov vortex core of a proximitised TI [15]. The search for Majorana zero-modes has been more advanced on one-dimensional (1D) semiconducting Rashba nanowires, proximity-coupled to a superconductor, and exposed to a strong magnetic field [7]; however, despite a remarkable experimental progress [16, 17], the distinction between topological and trivial ZBCPs remains contentious in this 1D platform [18, 19, 20]. It was recognized early on [21] and emphasized in the recent literature [22, 23, 24, 25] that a direct observation of the closing and reopening of the superconducting (SC) gap, which implies the topological phase-transition, is a more robust signature of topological superconductivity and thus Majorana zero-modes.
Figure 1: Phase-dependent Andreev bound states in a TI Josephson junction. a,b, Phase-dispersions of the perpendicular ($k_{y}$ = 0) mode in a TI line junction (a) and conventional Andreev bound states (ABS) with $\tau=0.9$ (b). In panel a, the red and blue lines correspond to the even- and odd-parity bound states, respectively. c, Evolution of the topological phase in the junction; here, the junction area is colored according to the gap function $m\propto\cos\frac{\varphi(y)}{2}$ (see the color scale). As $\varphi$ is tuned, the nature of the gap changes from trivial ($\varphi$ = 0, red) to topological ($\varphi$ = $2\pi$, blue). This is accompanied by a change in ground state parity from even to odd, as indicated by the extra electron (yellow) in the junction. In-between, the gap has to close ($\varphi=\pi$) via a gapless Majorana mode (green line). As the phase difference reaches $\varphi=4\pi$, a full period is completed. d, False-color SEM image of a typical device. The exfoliated TI flake (red) is contacted by Al (blue) forming an RF-SQUID geometry. Metallic contacts (Au) allow for probing the LDOS at two positions (TJ1 and TJ2) in the junction. e, color mapping of $dI/dV$ in the $B$ vs $V_{\rm b}$ plane measured on TJ1 at 50 mK with $V_{\rm G}$ = 0 V. At $\varphi$ = 0 (mod $2\pi$), a proximity-induced gap opens with conductance peaks at $\pm$130 $\mu$V. The gap closes and a broad ZBCP is observed at $\varphi=\pi$. A periodicity of 230 $\mu$T is observed, consistent with the device geometry. f, Line-cuts of the data in e at three selected magnetic fields corresponding to $\varphi$ = 0, $\frac{3}{4}\pi$, and $\pi$.
In this work, we turn to the TI platform and report the observation of robust gap closing and reopening in a TI Josephson junction upon changing the phase difference $\varphi$ across the junction. The spin-momentum locking of the TSS prohibits exact back-scattering, and hence the normal reflection is forbidden for a perpendicular mode at the boundary to a superconductor. Such an absence of normal back-scattering in turn guarantees perfect Andreev reflection, leading to a pair of perfectly transmitted, non-degenerate Andreev bound states (ABSs) in a short Josephson junction. These perfectly transmitted modes become gapless for $\varphi=\pi$, and the spin-non-degeneracy of the TSS makes the electron and hole branches to have opposite parity [26]. As $\varphi$ is varied, these nondegenerate modes cross and the ground-state parity alternates at $\varphi=(2n+1)\pi$ with $n$ integer [8]. Since the fermion parity is the $\mathbb{Z}_{2}$ topological invariant of a 1D superconductor [2], the gap closing and reopening corresponds to a topological phase transition. It is important to emphasize that the appearance of the perfectly transmitted ABSs is a direct consequence of the spin-momentum locking of the TSS that prohibits back-scattering. Therefore, as long as the Josephson effect takes place in the TSS, the topological phase transition as a function of $\varphi$ is a necessity. Nevertheless, this fundamental property has not been experimentally confirmed so far in TI Josephson junctions.
Note that the Fu-Kane theory is only valid for a constant $\varphi$ along the junction. When an out-of-plane magnetic field is applied to the junction, rendering $\varphi$ position-dependent [27], the gapless 1D mode becomes confined at the core of a Josephson vortex, where the local phase difference is $\varphi=(2n+1)\pi$. Because of this, the gapless 1D mode evolves into a Majorana zero-mode that is bound to the point defect formed by the Josephson vortex.
In our experiment, we observe periodic closings and reopenings of the gap by tuning $\varphi$ across the junction in a SQUID geometry and simultaneously detecting the local density of states (LDOS) at different locations. The gap closing, which is confirmed to be robust against a wide range of chemical potential, is accompanied by a multitude of low-lying excited states, as expected from theory. We further elucidate the locality of the involved states created by the phase gradient in perpendicular magnetic fields. Hence, our observations confirm the consequence of the spin-momentum locking in the TSS in a most direct manner. Based on this promising result, we suggest future experiments to nail down the topological superconductivity and to generate well-protected Majorana zero-modes.
Figure 2: Theoretical simulations. a, Color-mapping of the simulated differential conductance $G$ in units of the conductance quantum $G_{0}=e^{2}/h$ as a function of $V_{b}$ and $\varphi$ for a temperature $T=\Delta/(30k_{\rm B})$ in the Fu-Kane regime with $\varphi$ constant along the junction. b, Simulated behavior of the differential conductance $G(V_{b})$ at $\varphi$ = 0, $\frac{3}{4}\pi$, and $\pi$, which correspond to those shown in Fig. 1f. Note that in the simulation only the surface states and a single transversal mode are taken into account which explains the softer-gap in the experimental results of Fig. 1.
II Phase-biased TI Josephon junction
@@
The perpendicular mode of the ABSs (i.e. zero longitudinal momentum $k_{y}$) appearing in a line junction made on the TSS is expected to present a $4\pi$-periodic $\varphi$-dependence as shown in Fig. 1a in the short-junction limit. As mentioned above, this is dictated by the spin-momentum locking which prohibits back-scattering. The corresponding evolution of the gap in the junction is depicted in Fig. 1c: The gap is trivial for $0\leq\varphi<\pi$. It closes at the topological phase-transition with $\varphi=\pi$ where the gapless 1D mode (green line) appears, before reopening in the topological regime, with $\pi<\varphi<3\pi$. In the topological regime the parity of the ground state is odd, meaning that the bound state below the chemical potential $\mu$ is filled with an unpaired electron (shown in yellow). Because of this, the junction returns to the initial state only at $\varphi=4\pi$. In contrast, for a conventional Andreev bound state, a $2\pi$-periodicity is expected, as depicted in Fig. 1b. Here, the dispersion is given by [28]
@@
with $\tau$ the transmittance of the state and $\Delta$ the induced SC gap in the normal metal beneath the SC electrodes. In such a scenario, the gap can close in the case of perfect transmission ($\tau$ = 1), but the ground-state parity remains even across the gap closing, as conventional ABSs are doubly degenerate.
[…]
@@
Normal electrodes made of Au are deposited after putting a thin layer of $\mathrm{Al_{2}O_{3}}$ as a tunnel barrier. Due to the native oxide of Al, its surface can be treated as an insulator and the overlap to the Au electrodes does not contribute to the transport. A back-gate voltage can be applied via the degenerately-doped silicon substrate. Therefore, this setup enables us to directly probe the LDOS of the TI Josephson junction and to test the Fu-Kane theory [8]. Note that when the perpendicular magnetic field is very weak such that the large loop is threaded by only a few flux quanta (and no Josephson vortex nucleates in the junction), the phase $\varphi$ is almost constant along the junction, a situation akin to the Fu-Kane model. The case at larger magnetic field, when Josephson vortices nucleate in the junction, is discussed later. We focus on the results obtained from a device similar to the one shown in Fig. 1d with two independent tunnel junctions TJ1 and TJ2 located at different positions on the same Josephson junction. Additional data obtained from a second, essentially identical device, is shown in the supplement to demonstrate the reproducibility of our results.
III Tunnel spectra and gap closing
@@
In Fig. 1e, we show the continuous evolution of the differential conductance $dI/dV$ of TJ1 with the applied DC bias voltage $V_{b}$ and the perpendicular magnetic field $B$ around zero. Line cuts for selected fields are shown in Fig. 1f. Without magnetic field applied, we observe sharp particle-hole symmetric peaks at the bias voltages of ±$\pm 130\,\mathrm{\mu eV}$ and a strong reduction in $dI/dV$ at lower bias. The constant $dI/dV$ at higher bias corresponds to a resistance of about $400\,\mathrm{k\Omega}$, indicating we are in the tunnelling regime, where the observed $dI/dV$ is proportional to the LDOS. The peak energy of $130\,\mathrm{\mu eV}$ is smaller than the SC gap of our Al-film, which can be calculated from the critical temperature to be $170\,\mathrm{\mu eV}$. This confirms that the LDOS of the proximitised TI surface, not the LDOS of the Al electrode, is being measured. The LDOS at low bias does not drop to zero but remains finite, which indicates additional in-gap states that are not fully proximitised, consistent with our previous result on similar TI Josephson junctions, where the average transparency was $\tau_{\rm av}\approx 0.8$ [37]. Such a soft gap has also been reported for comparable devices based on graphene [31], carbon nanotube [33], HgTe [34], or InAs [35].
@@
As a finite phase bias $\varphi$ is generated across the junction with the application of an out-of-plane magnetic field, the positions of the particle-hole symmetric peaks shift to a lower bias voltage and their amplitude is diminished. For the magnetic field of about $115\,\mathrm{\mu T}$ that corresponds to $\varphi=\pi$, the particle-hole symmetric peaks merge into a single broad peak at zero bias, signalling the closing of the gap. The spectrum then evolves periodically, with the conductance at $230\,\mathrm{\mu T}$ becoming identical to the one observed at zero field. Having the Fu-Kane theory in mind, our data is consistent with the interpretation that the peak positions trace the gap, corresponding to the lowest-lying state at zero momentum, that evolves with $\varphi$ and closes at $\pi$ (see Fig. 1a). Above this gap (but still below the bulk gap), there is a continuum of states corresponding to finite momentum $k_{y}$ along the junction [8], which is indeed reflected in our LDOS spectra. Within this interpretation, the gap closing at $\varphi=\pi$ corresponds to a topological phase-transition [8]. It is useful to note that similar devices made on various materials in the ballistic regime [33, 34, 35] with $\tau<1$ exhibited gapped spectra at $\varphi=\pi$, consistent with Eq. (1). The gap closing observed here is most likely due to the unique property of the TSS, which necessitates the existence of a $\tau=1$ mode.
@@
To further corroborate this interpretation, we have performed numerical simulations of the differential conductance $G$ expected for the parameters of our devices (see supplement for details). Figure 2a shows the calculated $G$ as a function of $V_{b}$ and $\varphi$, while the line cuts for the same $\varphi$ values as in Fig. 1f are shown in Fig. 2b. These simulations confirm that the even- and odd-parity bound states having a $4\pi$-periodicity produce the two branches that cross at $\varphi=(2n+1)\pi$, as observed in experiments. In Fig. 2b, the coherence peaks show up at the edge of the excitation gap and they diminish rapidly with increasing $\varphi$. Our simulations suggest that the exact shape of the spectrum at $\varphi=\pi$ depends sensitively on the experimental conditions, and the gap closing can result in a nearly flat $dI/dV$ (see supplement). All these results are consistent with our experimental observations. It should be emphasized that the gap closing always occurs at $\varphi=(2n+1)\pi$, but it does not need to be accompanied by a ZBCP, according to our simulations.
IV Locality of the gap closing in perpendicular magnetic fields
Figure 3: Locality of the gap closing. a, Schematic drawing of a planar TI Josephson junction in a perpendicular magnetic field, which generates a position-dependent phase $\varphi(y)$ across the junction. The distribution of the topological gap function $m\propto\cos\frac{\varphi(y)}{2}$ is depicted in the junction with a color scale. The position of $\varphi(y)=(2n+1)\pi$ corresponds to the Josephson vortex-core, where the gap closes (white color) and a Majorana zero-mode (green dot) appears. b, Position-dependent gap $|m|=|\cos\frac{\varphi(y)}{2}|$ near a Josephson vortex-core in a magnetic field of $B$ = 5 mT, plotted with the excited states calculated from the model of Ref. 27 for a chemical potential $\mu$ = 25 meV and $\Delta$ = 0.17 meV (thin horizontal lines), which form a quasi-continuum with a minigap of only 2% of $\Delta$. The vertical red lines mark the spatial extent of the expected Majorana zero-mode appearing around $y$ = 0. c, color mapping of $dI/dV$ in the $B$ vs $V_{\rm b}$ plane for TJ1 (top panel) and TJ2 (bottom panel) measured in the magnetic field range of 5 – 6 mT at 50 mK with $V_{\rm G}=-5$ V. Note the slight difference in the periodicity, which is 230 $\mu$T for TJ1 and 200 $\mu$T for TJ2. The middle panel depicts how the distribution of the topological gap function in the junction calculated with Eq. (2) evolves with magnetic fields; the position where the gap function locally becomes zero with $\varphi(y)=(2n+1)\pi$ corresponds to local gap closing (white regions).